# EMFusion: Conditional Diffusion Framework for Trustworthy Frequency Selective EMF Forecasting in Wireless Networks

### v1

@@ III-A Multivariate Conditional Diffusion Model
We consider a multivariate time-series forecasting problem in which each observation $\mathbf{x}_{t}\in\mathbb{R}^{N}$ represents measurements from $N$ variates or frequency channels at time $t$. Given a historical window of length $H$, denoted by $\mathbf{X}_{\text{past}}=(\mathbf{x}_{T-H+1},\ldots,\mathbf{x}_{T})$, the objective is to predict the future multichannel trajectory $\mathbf{X}_{\text{future}}=(\mathbf{x}_{T+1},\ldots,\mathbf{x}_{T+F})$ over a forecast horizon $F$. To model the conditional distribution $p_{\theta}(\mathbf{X}_{\text{future}}\mid\mathbf{X}_{\text{past}},\mathbf{c})$, we introduce an external condition vector $\mathbf{c}$, which may include calendar features, operational metadata, or other exogenous covariates. Unlike vectorized formulations, we treat the full future window as a matrix $\mathbf{X}_{0}:=[\mathbf{X}_{\text{past}},\mathbf{X}_{\text{future}}]\in\mathbb{R}^{(H+F)\times N}$ and apply diffusion directly in this multivariate tensor space. This enables probabilistic forecasting while preserving the temporal–channel structure of the signal.
In the diffusion-based formulation, the forward noising process corrupts the clean $\mathbf{X}_{0}$ over $T$ time steps according to the Markov transition process given by [ho2020ddpm, huang2025probabilistic]:
@@
and $\mathbf{c}$ encapsulates features from the dataset, such as Italy_WorkingDay, Italy_WorkingHour, and Italy_Season, which depict factors influencing EMF variations [meijer2024rise, su2025diffusion]. By conditioning the noise prediction on $\mathbf{c}$, the model learns to generate EMF trajectories consistent with temporal continuity and relevant external dependencies.
@@
During inference, a classifier-free guidance (CFG) mechanism [ref:classifier-free, huang2025probabilistic] steers the generation process. The final noise prediction is obtained by combining the conditional and unconditional outputs of the denoising network:
@@
TABLE II: Frequency $f$ (MHz) allocation for mobile operators and technology band used in the Italian dataset. Each entry shows the antenna factor $(\alpha_{f})$ in dB m-1 from Keysight N6850A calibration [keysight_n6850a].
@@ III-B1 Residual Block with Time-Step Embedding
The core of the $\epsilon_{\theta}$ network is a stack of residual blocks that progressively refine the feature maps. As shown in Fig. 1, each residual block applies normalization, nonlinearity, convolution, and time step conditioning, followed by a residual shortcut. Given an input feature map $\mathbf{h}$ from the previous layer, the block first applies Group Normalization and a Sigmoid Linear Unit (SiLU) activation, followed by a $3\times 3$ convolution, i.e., [ronneberger2015u]:
@@
The updated feature map $\hat{\mathbf{h}}$ is then processed by a second sequence of operations, GroupNorm, SiLU, dropout, and another $3\times 3$ convolution [ronneberger2015u]:
@@
Finally, a residual shortcut is applied as [he2016deep]:
@@ III-B2 Cross-Attention Module
To compute the attention, the 3D feature map $\mathbf{h}$ is first flattened into a 2D tensor, $\mathbf{h}_{\text{flat}}$, with shape $(C_{h},N_{h})$, where $N_{h}=h_{h}\cdot w_{h}$. Since cross-attention calculation basically contains linear connection without recurrence, it’s important to fully make use of the order of the sequence. Positional encoding adds positional information to $\mathbf{c}$ and $\mathbf{h}_{\text{flat}}$, which helps the module extract temporal and spatial correlations. Consequently, we add positional encoding to $\mathbf{c}$ and $\mathbf{h}_{\text{flat}}$ according to [ref:attention]:
@@
where $W_{Q}$, $W_{K}$, and $W_{V}$ are learnable weight matrices [ref:latent, huang2025probabilistic]. The attention matrix $A$ and the final output are calculated using the standard scaled dot-product attention mechanism:
@@
where $e_{k}$ is the dimension of the keys (and queries), used for scaling. According to ① in Fig. 2, a single row of the attention matrix $A$ represents the set of attention weights that one specific element from the flattened feature map $\mathbf{h}_{\text{flat}}$ (represented by a row of $Q$) pays to all $N_{f}$ elements of the condition vector $\mathbf{c}$ (represented by the rows of $K$). ② in Fig. 2 shows a column of the value $V$ that corresponds to a feature map weighted by the attention in $A$ [petit2021u]. The output of this operation, $AV$, represents the contextually relevant information from $\mathbf{c}$, weighted according to its relevance to $\mathbf{h}$. This output is then reshaped and projected to match the original feature map’s dimensions ($C_{h}\times h_{h}\times w_{h}$). This result is finally added back to the original feature map $\mathbf{h}$ via a residual connection, which is shown in the previous section:
@@ III-E Probabilistic Interval Construction
Subsection III-E describes probabilistic interval construction for calibrated uncertainty quantification. The procedure begins once the EMFusion has generated an ensemble of $\hat{N}$ distinct future EMF trajectories $\mathbf{S}=\{\hat{\mathbf{X}}_{\mathrm{future}}^{i}\}_{i=1,\cdots,\hat{N}}\in\mathbb{R}^{\hat{N}\times F}$ for each of the $N$ monitored carriers (e.g., individual 2G, 3G, 4G, and 5G carriers across multiple operators). For each future time step $k\in\{1,2,\cdots,F\}$, this yields a set of $\hat{N}$ predicted values, denoted as $\mathbf{Z}=\{z_{1},z_{2},\dots,z_{\hat{N}}\}$. This ensemble can be viewed as a discrete, non-parametric sample drawn from the learned conditional predictive distribution. To convert this sample set into a continuous probability density function (PDF), Kernel Density Estimation (KDE) is employed [huang2025probabilistic]. KDE is a classical non-parametric technique that estimates the underlying density $f(z)$ by placing a kernel function, typically Gaussian, at each sample point and summing the contributions:
$f(z)=\dfrac{1}{\hat{N}\cdot s_{p}}\sum\limits_{i=1}^{\hat{N}}K\left(\dfrac{z-z_{i}}{s_{p}}\right),$ (22)
where $K(\cdot)$ is the kernel function and $s_{p}$ is the bandwidth parameter. The bandwidth $s_{p}$ is crucial for controlling smoothness, determining the bias and variance trade-off in the estimated density. This nonparametric statistical method can directly estimate probability distributions from discrete scenarios without requiring prior assumptions about the distribution. Once the PDF $f(z)$ is derived, the cumulative distribution function $F(z)$ is computed, and the prediction interval is obtained via quantile extraction. Specifically, for a target confidence level $\gamma$, the lower and upper quantile bounds $q_{\underline{\alpha}}$ and $q_{\overline{\alpha}}$ satisfy $PI_{\gamma}=[q_{\underline{\alpha}},q_{\overline{\alpha}}]$, $\qquad q_{\alpha}=F^{-1}(\alpha)$, where $\gamma=\overline{\alpha}-\underline{\alpha}$, and $\underline{\alpha}$, $\overline{\alpha}$ denote the lower and upper quantile levels, respectively.
In most KDE-based interval estimation scenarios, symmetric intervals satisfying $\overline{\alpha}+\underline{\alpha}=1$ are adopted. In this work, however, since EMFusion directly generate multiple stochastic trajectories, we employ a simplified empirical interval construction approach. Specifically, given sorted samples $z_{1}<z_{2}<\dots<z_{\hat{N}}$, the interval $PI_{\gamma}$ is directly determined from the corresponding order statistics without KDE, offering a computationally efficient alternative:
$\hat{N}_{\gamma}=\lfloor(1-\gamma)\cdot\hat{N}\rfloor$ (23)
$PI_{\gamma}=[z_{{\hat{N}}_{\gamma}},z_{\hat{N}-\hat{N}_{\gamma}}]$ (24)
@@ IV-A1 EMF Monitoring Set-up
The EMF monitoring chain couples an omnidirectional antenna, an Anritsu MS27102A remote spectrum-monitoring unit, and the ARPA-Lazio Search24 control software [chiaraviglio2022six]. The Keysight N6850A passive antenna [keysight_n6850a] is connected to the portable spectrum analyzer via a 1.5 m low-loss coaxial cable. The antenna delivers uniform gain across all licensed cellular bands up to the 3.6 GHz. The weather-proof MS27102A sweeps 9 kHz–6 GHz continuously and, thanks to its internal pre-selector, outputs real-time narrow-band channel power while operating unattended on a rooftop site [ms27102a]. Search24 [search24] scripts enable streaming the data to a local computer over a dedicated Ethernet link.
@@ IV-A2 Conversion of Received Power to EMF
The received power for a given frequency channel $f$ at the output of the antenna is measured in dBm. For a given frequency $f$, the received power is then converted into the electric-field strength (measured in V/m) as given below [chiaraviglio2021health]:
@@
where $c$ is the speed of light. $A_{GC}$ is antenna-gain calibration constant. The antenna factor $\alpha_{f}$ changes as a function of frequency $f$ as in [chiaraviglio2022six, keysight_n6850a_antfac].
@@ IV-A3 External Conditions
  • Italy_WorkingDay $\in\{0,1\}$: indicates whether $t$ falls on a weekday (Monday–Friday) and is not listed in the Italian public–holiday calendar. [^2] The holiday list covers national and widely observed regional holidays for 2023–2024, including Easter, Liberation Day, Labour Day, Republic Day, Ferragosto, All Saints’ Day, Immaculate Conception, and the Christmas–New Year period. It also includes the university closure date and no teaching activities dates for Università degli Studi di Roma Tor Vergata.[TorVergataAcademicCalendarPage2025]
@@ IV-B EMF Classification and Correlation Analysis
TABLE IV: A comparative analysis of univariate forecasting performance of EMFusion with imputation without external conditions. Metrics are ordered as CRPS, ND, NRMSE, MAPE, and PICP. Lower is better for CRPS, ND, NRMSE, MAPE; higher is better for PICP. Bold = best in row; underline = second best in row.
@@ V-A2 Probabilistic Forecast Metrics
$\bullet$ Prediction Interval Coverage Probability (PICP): measures the percentage of true values that fall within the model’s predicted interval. Let the prediction interval at time step $i$ be defined by its lower and upper quantiles, $L_{i}=q_{\underline{\alpha}}^{\,i}$ and $U_{i}=q_{\overline{\alpha}}^{\,i}$, corresponding to a nominal confidence level $\gamma=1-\alpha$. For example, an 80% prediction interval corresponds to $\gamma=0.8$ and $\alpha=0.2$. We define the coverage indicator as follows:
@@
The PICP is then computed as follows:
$\mathrm{PICP}=\frac{100}{m}\sum_{i=1}^{m}c_{i}.$ (33)
Ideally, $\mathrm{PICP}$ should be close to the nominal value $100\gamma$, indicating well-calibrated predictive intervals.
@@
and compute CRPS at each time step accordingly [RasulEtAl2021TimeGrad].
@@ V-B Considered Baselines
To validate the performance of EMFusion, we conduct a comparative analysis against a suite of forecasting models. A brief description of each basline model is provided below:
  • DDPM [ref:diffusion-energy]: the model uses a U-Net denoiser with time-dependent noise embeddings and is trained to match the true forward diffusion process via noise-prediction.
  • WGAN [huang2025probabilistic]: the generator maps historical EMF trajectories and latent noise to future EMF scenarios, while the critic enforces distributional realism through the Wasserstein loss with gradient penalty.
  • VAE [ref:vae]: the encoder maps historical EMF sequences into a low-dimensional latent space, and the decoder reconstructs future EMF trajectories from sampled latent variables conditioned on the history. This yields a parametric predictive distribution via the latent Gaussian prior.
  • Normalizing Flow (NF) [ref:diffusion-energy]: Starting from a base distribution (e.g., multivariate Gaussian), a sequence of invertible transformations is learned to model the joint distribution of future EMF values given the past values, enabling exact likelihood evaluation and sampling.
  • Improved Quantile LSTM (IQLSTM): follows [ref:iqlstm], where an LSTM-based temporal encoder outputs multiple conditional quantiles of the future EMF distribution. The model is trained using a pinball (quantile) loss over a set of pre-defined quantile levels, providing a non-parametric representation of forecast uncertainty for each time step.
  • Dropout-based Probabilistic Model: Following the Monte Carlo Dropout approach of Gal and Ghahramani [ref:dropout], this algorithm employs a deep LSTM forecaster with dropout layers activated both during training and inference, byy running multiple stochastic forward passes with different dropout masks
  • TimeGrad: TimeGrad [RasulEtAl2021TimeGrad] is a diffusion-based time-series forecasting model that combines an autoregressive RNN backbone with a conditional diffusion head.

### v2

@@ III-A Multivariate Conditional Diffusion Model
We consider a multivariate time-series forecasting problem in which each observation $\mathbf{x}_{k}\in\mathbb{R}^{N}$ represents measurements from $N$ variates or frequency channels at time $k$. Given a historical window of length $H$, denoted by $\mathbf{X}_{\text{past}}=(\mathbf{x}_{-(H+1)},\ldots,\mathbf{x}_{0})$, the objective is to predict the future multichannel trajectory $\mathbf{X}_{\text{future}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{F})$ over a forecast horizon $F$. To model the conditional distribution $p_{\theta}(\mathbf{X}_{\text{future}}\mid\mathbf{X}_{\text{past}},\mathbf{c})$, we introduce an external condition vector $\mathbf{c}$, which may include calendar features, operational metadata, or other exogenous covariates. Unlike vectorized formulations, we treat the full future window as a matrix $\mathbf{X}_{0}:=[\mathbf{X}_{\text{past}},\mathbf{X}_{\text{future}}]\in\mathbb{R}^{(H+F)\times N}$ and apply diffusion directly in this multivariate tensor space. This enables probabilistic forecasting while preserving the temporal–channel structure of the signal.
In the diffusion-based formulation, the forward noising process corrupts the clean $\mathbf{X}_{0}$ over $T$ time steps according to the Markov transition process given by [60, 23]:
@@
and $\mathbf{c}$ encapsulates features from the dataset, such as Italy_WorkingDay, Italy_WorkingHour, and Italy_Season, which depict factors influencing EMF variations [58, 59]. By conditioning the noise prediction on $\mathbf{c}$, the model learns to generate EMF trajectories consistent with temporal continuity and relevant external dependencies.
@@
During inference, a classifier-free guidance (CFG) mechanism [61, 23] steers the generation process. The final noise prediction is obtained by combining the conditional and unconditional outputs of the denoising network:
@@
TABLE II: Frequency $f$ (MHz) allocation for mobile operators and technology band used in the Italian dataset. Each entry shows the antenna factor $(\alpha_{f})$ in dB m-1 from Keysight N6850A calibration [62].
@@ III-B1 Residual Block with Time-Step Embedding
The core of the $\epsilon_{\theta}$ network is a stack of residual blocks that progressively refine the feature maps. As shown in Fig. 1, each residual block applies normalization, nonlinearity, convolution, and time step conditioning, followed by a residual shortcut. Given an input feature map $\mathbf{h}$ from the previous layer, the block first applies Group Normalization and a Sigmoid Linear Unit (SiLU) activation, followed by a $3\times 3$ convolution, i.e., [63]:
@@
The updated feature map $\hat{\mathbf{h}}$ is then processed by a second sequence of operations, GroupNorm, SiLU, dropout, and another $3\times 3$ convolution [63]:
@@
Finally, a residual shortcut is applied as [64]:
@@ III-B2 Cross-Attention Module
To compute the attention, the 3D feature map $\mathbf{h}$ is first flattened into a 2D tensor, $\mathbf{h}_{\text{flat}}$, with shape $(C_{h},N_{h})$, where $N_{h}=h_{h}\cdot w_{h}$. Since cross-attention calculation basically contains linear connection without recurrence, it’s important to fully make use of the order of the sequence. Positional encoding adds positional information to $\mathbf{c}$ and $\mathbf{h}_{\text{flat}}$, which helps the module extract temporal and spatial correlations. Consequently, we add positional encoding to $\mathbf{c}$ and $\mathbf{h}_{\text{flat}}$ according to [65]:
@@
where $W_{Q}$, $W_{K}$, and $W_{V}$ are learnable weight matrices [66, 23]. The attention matrix $A$ and the final output are calculated using the standard scaled dot-product attention mechanism:
@@
where $e_{k}$ is the dimension of the keys (and queries), used for scaling. According to ① in Fig. 2, a single row of the attention matrix $A$ represents the set of attention weights that one specific element from the flattened feature map $\mathbf{h}_{\text{flat}}$ (represented by a row of $Q$) pays to all $N_{f}$ elements of the condition vector $\mathbf{c}$ (represented by the rows of $K$). ② in Fig. 2 shows a column of the value $V$ that corresponds to a feature map weighted by the attention in $A$ [67]. The output of this operation, $AV$, represents the contextually relevant information from $\mathbf{c}$, weighted according to its relevance to $\mathbf{h}$. This output is then reshaped and projected to match the original feature map’s dimensions ($C_{h}\times h_{h}\times w_{h}$). This result is finally added back to the original feature map $\mathbf{h}$ via a residual connection, which is shown in the previous section:
@@ III-E Probabilistic Interval Construction
Once EMFusion generates an ensemble of $\hat{N}$ future EMF trajectories $\mathbf{S}=\{\hat{\mathbf{X}}_{\mathrm{future}}^{i}\}_{i=1}^{\hat{N}}\in\mathbb{R}^{\hat{N}\times F}$ for each monitored carrier (e.g., individual 2G, 3G, 4G, and 5G carriers across multiple operators), the samples at forecast horizon $k\in\{1,\dots,F\}$ are given by $\mathbf{Z}_{k}=\{z_{k}^{1},\dots,z_{k}^{\hat{N}}\}$. These samples are interpreted as Monte-Carlo draws from the learned conditional predictive distribution. A smooth estimate of the predictive density at horizon $k$ is obtained via Kernel Density Estimation (KDE) as shown below:
$\hat{f}_{k}(z)=\frac{1}{\hat{N}s_{p}}\sum_{i=1}^{\hat{N}}K\!\left(\frac{z-z_{k}^{i}}{s_{p}}\right),$ (22)
where $K(\cdot)$ denotes the Gaussian kernel function and $s_{p}$ is the bandwidth parameter controlling the bias–variance trade-off. The corresponding cumulative distribution function is $\hat{F}_{k}(z)=\int_{-\infty}^{z}\hat{f}_{k}(t)\,dt.$ For a target confidence level $\gamma$, define $\underline{\alpha}=\frac{1-\gamma}{2},\qquad\overline{\alpha}=1-\underline{\alpha}.$ The prediction interval based on the estimated density is given by:
$PI_{\gamma}=\left[\hat{F}_{k}^{-1}(\underline{\alpha}),\hat{F}_{k}^{-1}(\overline{\alpha})\right].$ (23)
In practical implementation, instead of explicitly evaluating the KDE and numerically inverting the CDF, the required quantiles are computed directly from the ordered ensemble samples. Let $z_{k}^{(1)}\leq\dots\leq z_{k}^{(\hat{N})}$ denotes the order statistics of $\mathbf{Z}_{k}$. Let us define $\ell=\left\lceil\underline{\alpha}\hat{N}\right\rceil,\quad u=\left\lceil\overline{\alpha}\hat{N}\right\rceil$. The empirical prediction interval is then given by $PI_{\gamma}=[L_{k},U_{k}]=\left[z_{k}^{(\ell)},z_{k}^{(u)}\right].$ As the ensemble size $\hat{N}$ increases, the empirical quantiles converge to those of the underlying distribution, ensuring statistically consistent and efficient uncertainty estimation.
@@ IV-A1 EMF Monitoring Set-up
The EMF monitoring chain couples an omnidirectional antenna, an Anritsu MS27102A remote spectrum-monitoring unit, and the ARPA-Lazio Search24 control software [68]. The Keysight N6850A passive antenna [62] is connected to the portable spectrum analyzer via a 1.5 m low-loss coaxial cable. The antenna delivers uniform gain across all licensed cellular bands up to the 3.6 GHz. The weather-proof MS27102A sweeps 9 kHz–6 GHz continuously and, thanks to its internal pre-selector, outputs real-time narrow-band channel power while operating unattended on a rooftop site [69]. Search24 [70] scripts enable streaming the data to a local computer over a dedicated Ethernet link.
@@ IV-A2 Conversion of Received Power to EMF
The received power for a given frequency channel $f$ at the output of the antenna is measured in dBm. For a given frequency $f$, the received power is then converted into the electric-field strength (measured in V/m) as given below [2]:
@@
where $c$ is the speed of light. $A_{GC}$ is antenna-gain calibration constant. The antenna factor $\alpha_{f}$ changes as a function of frequency $f$ as in [68, 71].
@@ IV-A3 External Conditions
  • Italy_WorkingDay $\in\{0,1\}$: indicates whether $t$ falls on a weekday (Monday–Friday) and is not listed in the Italian public–holiday calendar. [^3] The holiday list covers national and widely observed regional holidays for 2023–2024, including Easter, Liberation Day, Labour Day, Republic Day, Ferragosto, All Saints’ Day, Immaculate Conception, and the Christmas–New Year period. It also includes the university closure date and no teaching activities dates for Università degli Studi di Roma Tor Vergata.[72]
@@ IV-B EMF Classification and Correlation Analysis
TABLE IV: Forecasting performance of EMFusion without external conditions. Metrics are ordered as CRPS, ND, NRMSE, MAPE, and PICP. Lower is better for CRPS, ND, NRMSE, MAPE; higher is better for PICP. Bold = best in row; underline = second best in row. PICP values lower than 10 are considered abnormal and replaced with “–”.
@@ V-A2 Probabilistic Forecast Metrics
$\bullet$ Prediction Interval Coverage Probability (PICP): measures the proportion of realized values that fall within the predicted interval. For each forecast horizon step $k$, let the prediction interval be $PI_{\gamma}=[L_{k},U_{k}],$ where $\underline{\alpha}=\frac{1-\gamma}{2},\qquad\overline{\alpha}=1-\underline{\alpha}.$ For example, an 80% prediction interval corresponds to $\gamma=0.8$, yielding $\underline{\alpha}=0.1$ and $\overline{\alpha}=0.9$. Ideally, $PI_{\gamma}$ should closely match the target confidence level $\gamma$. Define the coverage indicator as follows:
@@
Over $m$ forecasting values, the PICP is computed as
$\mathrm{PICP}=\frac{1}{m}\sum_{i=1}^{m}c_{i}.$ (32)
For reporting in percentage form, $100\times\mathrm{PICP}$ is used. Ideally, $\mathrm{PICP}\approx\gamma$, indicating well-calibrated predictive intervals.
@@
and compute CRPS at each time step accordingly [73].
@@ V-B Considered Baselines
To validate the performance of EMFusion, we conduct a comparative analysis against a suite of forecasting models. A brief description [^4] The considered baselines are inherently unconditional and do not incorporate external contextual information. For fair comparison, we consider unconditional version of EMFusion (Table III) under identical contextual settings. All EMFusion variants in Tables IV and V share the same backbone architecture, parameter count, optimization settings, and training protocol. of each basline model is provided below:
  • DDPM [74]: the model uses a U-Net denoiser with time-dependent noise embeddings and is trained to match the true forward diffusion process via noise-prediction.
  • WGAN [23]: the generator maps historical EMF trajectories and latent noise to future EMF scenarios, while the critic enforces distributional realism through the Wasserstein loss with gradient penalty.
  • VAE [75]: the encoder maps historical EMF sequences into a low-dimensional latent space, and the decoder reconstructs future EMF trajectories from sampled latent variables conditioned on the history. This yields a parametric predictive distribution via the latent Gaussian prior.
  • Normalizing Flow (NF) [74]: Starting from a base distribution (e.g., multivariate Gaussian), a sequence of invertible transformations is learned to model the joint distribution of future EMF values given the past values, enabling exact likelihood evaluation and sampling.
  • Improved Quantile LSTM (IQLSTM): follows [76], where an LSTM-based temporal encoder outputs multiple conditional quantiles of the future EMF distribution. The model is trained using a pinball (quantile) loss over a set of pre-defined quantile levels, providing a non-parametric representation of forecast uncertainty for each time step.
  • Dropout-based Probabilistic Model: Following the Monte Carlo Dropout approach of Gal and Ghahramani [77], this algorithm employs a deep LSTM forecaster with dropout layers activated both during training and inference, byy running multiple stochastic forward passes with different dropout masks
  • TimeGrad: TimeGrad [73] is a diffusion-based time-series forecasting model that combines an autoregressive RNN backbone with a conditional diffusion head.